# The Angular Momentum Penrose Inequality

### v2

@@ 7 Stage 4: Sub-Extremality
Step 3: Preservation along AMO flow. The Dain–Reiris inequality $A(\Sigma)\geq 8\pi|J(\Sigma)|$ is established in [22] using variational methods specific to MOTS. We do not re-derive this inequality here; instead, we show that once it holds at $t=0$, it is preserved along the AMO flow by the following simple argument:
@@
    AA is non-decreasing: By the AMO area monotonicity (which requires only $R_{\tilde{g}}\geq 0$, established in Theorem 5.6), we have $A^{\prime}(t)\geq 0$ for almost all $t$.
@@
    Dain–Reiris inequality (Theorem 7.1): The proof of $A\geq 8\pi|J|$ in [22] crucially uses the stability condition through a variational argument. Unstable MOTS can violate this bound.
@@ 8 Synthesis: Complete Proof
Hypothesis Usage Summary. The four hypotheses enter the proof as follows (Stage references point to major sections of the paper; Step references are to the proof below): (H1) DEC: Ensures $R_{\tilde{g}}\geq 0$ after conformal transformation (Stage 2/Step 2), which drives AMO monotonicity (Step 6). (H2) Axisymmetry: Defines angular momentum $J$, guarantees axisymmetric solutions at every stage, and enables $J$-conservation (Stage 3/Step 4). (H3) Exterior vacuum: Ensures Komar and ADM angular momenta coincide; enables clean asymptotics for boundary evaluation (Step 7). (H4) Strictly stable MOTS: Guarantees $|\theta^{-}|>0$, ensuring proper cylindrical blow-up and correct boundary values at $t=0$ (Step 7).
@@
Let $(M,g,K)$ be asymptotically flat, axisymmetric data satisfying DEC with outermost stable MOTS $\Sigma$. The proof proceeds in seven steps, where Steps 1–4 correspond to the major Stages 1–4 of the paper.
Step 1 (Jang equation, Stage 1): By Theorem 4.12, solve the axisymmetric Jang equation to obtain $(\bar{M},\bar{g})$ with cylindrical ends at $\Sigma$.
Step 2 (Lichnerowicz equation, Stage 2): By Theorem 5.6, solve the AM-Lichnerowicz equation to obtain $\tilde{g}=\phi^{4}\bar{g}$ with $R_{\tilde{g}}\geq 0$.
@@
Step 4 ($J$-conservation, Stage 3): By Theorem 6.10, $J(t)=J$ for all $t\in[0,1]$.
Step 5 (Sub-extremality, Stage 4): By Theorem 7.1, $A(t)\geq 8\pi|J|$ for all $t$.
Step 6 (Monotonicity): By Theorem 6.22, $m_{H,J}(t)$ is monotone increasing.
Step 7 (Boundary values): Boundary values as $p\to 1^{+}$.
@@
Step 1: Geometric setup on the Jang manifold. On the Jang manifold $(\bar{M},\bar{g})$, the MOTS $\Sigma$ becomes the boundary of the cylindrical end. The key property is that the mean curvature $H_{\bar{g}}$ of $\Sigma$ in $(\bar{M},\bar{g})$ vanishes, i.e., $\Sigma$ is a minimal surface in the Jang metric. We now prove this crucial fact.
Detailed derivation of $H_{\bar{g}}|_{\Sigma}=0$: The Jang surface $\Gamma_{f}=\{(x,f(x)):x\in M\}$ is embedded in $(M\times\mathbb{R},g+dt^{2})$. Near the MOTS $\Sigma$, the Jang solution $f$ blows up as:
@@
$\bar{g}=g+df\otimes df=g+\frac{ds\otimes ds}{s^{2}}+O(1).$
In the cylindrical coordinate $t=-\ln s$ (so $s=e^{-t}$ and $ds=-e^{-t}dt$), this becomes:
@@
Now, the mean curvature of $\Sigma_{t}:=\{t\}\times\Sigma$ in the exact cylinder $\mathbb{R}_{+}\times\Sigma$ with product metric is zero, since $\Sigma_{t}$ are totally geodesic slices. In the actual Jang metric $\bar{g}$, the correction $O(e^{-\beta_{0}t})$ contributes:
$H_{\bar{g}}(\Sigma_{t})=O(e^{-\beta_{0}t})\to 0\quad\text{as }t\to\infty.$
@@
$H_{\bar{g}}|_{\Sigma}:=\lim_{t\to\infty}H_{\bar{g}}(\Sigma_{t})=0.$
@@
This argument can also be seen variationally: the AM-Lichnerowicz equation is the Euler–Lagrange equation for the functional $\mathcal{E}[\phi]=\int 8|\nabla\phi|^{2}+R_{\bar{g}}\phi^{2}+\frac{\Lambda_{J}}{6}\phi^{-6}$. On a manifold with minimal boundary, the natural boundary condition for critical points is Neumann: $\partial_{\nu}\phi=0$ (see [41, Proposition 3.2]).
@@
$f(s,y)=C_{0}\ln(1/s)+A(y)+O(s^{\alpha})=C_{0}t+A(y)+O(e^{-\alpha t}).$
@@
    Limit of $m_{H,J}(t)$: The value $m_{H,J}(0)$ is defined as $\lim_{t\to 0^{+}}m_{H,J}(t)$. By the continuity of area and the fact that $\Sigma_{t}\to\Sigma$ in the Hausdorff topology (with controlled curvature from the $p$-harmonic structure), this limit equals the AM-Hawking mass computed directly on $\Sigma$ via Lemmas 8.1 and 8.2.
@@
    Mean curvature regularity: Since the level sets $\Sigma_{t}$ are $C^{1,\alpha}$ regular for $p\in(1,2]$ [1], the mean curvature $H$ and second fundamental form $h$ are well-defined almost everywhere. The Hawking mass integral $\int_{\Sigma_{t}}H^{2}\,dA$ is finite for regular level sets.
@@
    $m_{H}(t)$ $=\sqrt{\frac{A(t)}{16\pi}}\left(1-W(t)\right)^{1/2}$
    $\approx\frac{r(t)}{2}\left(\frac{2M_{\mathrm{ADM}}}{r(t)}+O(r(t)^{-1-\tau})\right)^{1/2}$
@@
    The expansion uses the standard ADM mass formula: for coordinate spheres $S_{r}$, $\int_{S_{r}}H^{2}dA=16\pi-32\pi M_{\mathrm{ADM}}/r+O(r^{-1-\tau})$. ∎
@@
Conclusion: By the monotonicity from Step 6 and the mass chain from Lemma 5.12:
@@
The last inequality uses the lower bound analysis from Step 7 at the MOTS, which becomes an equality for Kerr initial data. ∎

### v3

@@ 7 Stage 4: Sub-Extremality
Remark 7.1 (Logical Ordering and Non-Circularity).
This section establishes the sub-extremality bound $A(t)\geq 8\pi|J|$ for AMO level sets. To prevent circularity, we explicitly state the logical dependencies:
What is already established at this stage:
  1. (1)
    Stage 1 (Section 4): The Jang manifold $(\bar{M},\bar{g})$ exists with cylindrical end at the MOTS $\Sigma$.
  2. (2)
[…]
  3. (3)
    Stage 3 (Section 6): The AMO $p$-harmonic potential $u_{p}$ exists, foliating $\tilde{M}$ by level sets $\{\Sigma_{t}\}_{t\in[0,1]}$.
  4. (4)
    Angular momentum conservation (Theorem 6.14 in Section 6): $J(\Sigma_{t})=J$ for all $t$.
What this section proves:
  • The initial sub-extremality $A(\Sigma)\geq 8\pi|J(\Sigma)|$ follows from the Dain–Reiris theorem [22], which depends only on the initial data $(M,g,K)$.
  • The preservation along the flow uses the $J$-conservation from Stage 4 and the area lower bound from the Dain–Reiris theorem applied to each $\Sigma_{t}$.
What this section does NOT use:
  • The sub-extremality bound does not use the AMO monotonicity (Stage 6).
  • The sub-extremality bound does not use the final inequality $m_{H,J}(0)\leq M_{\mathrm{ADM}}$.
This ordering ensures no circular reasoning: sub-extremality is an input to the mass monotonicity formula, not an output.
@@
Step 3: Preservation along AMO flow. The Dain–Reiris inequality $A(\Sigma)\geq 8\pi|J(\Sigma)|$ is established in [22] using variational methods specific to MOTS. We do not re-derive this inequality here; instead, we show that once it holds at $t=0$, it is preserved along the AMO flow by the following rigorous argument:
@@
    AA is non-decreasing: By the AMO area monotonicity, we establish that $A^{\prime}(t)\geq 0$ for almost all $t\in(0,1)$. We provide a complete proof:
[…]
    $A^{\prime}(t)=\int_{\Sigma_{t}}\frac{1}{|\nabla u|}\left(R_{\tilde{g}}+2|\mathring{h}|^{2}+\frac{2}{(p-1)^{2}}\left(H-(p-1)\frac{\Delta u}{|\nabla u|}\right)^{2}\right)d\sigma.$ (92)
    Each term in the integrand is non-negative:
    • $R_{\tilde{g}}\geq 0$ by Theorem 5.8 (AM-Lichnerowicz equation);
    • $|\mathring{h}|^{2}\geq 0$ (squared norm of traceless second fundamental form);
    • The third term is a squared quantity, hence non-negative.
    Since $|\nabla u|>0$ on regular level sets (which comprise all but a measure-zero set of $t$ values by Sard’s theorem), we conclude $A^{\prime}(t)\geq 0$ for a.e. $t\in(0,1)$.
    Remarks on the derivation:
    1. (a)
      The AMO formula (92) is derived using the Bochner identity, the $p$-harmonic equation, and integration by parts—see [1, Theorem 3.1] or the self-contained derivation in Proposition 6.25.
    2. (b)
      The condition $R_{\tilde{g}}\geq 0$ is essential: the conformal transformation $\tilde{g}=\phi^{4}\bar{g}$ with $\phi$ solving the AM-Lichnerowicz equation ensures $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$. Without this, $R_{\tilde{g}}$ could be negative and area monotonicity would fail.
    3. (c)
      For the limit $p\to 1^{+}$, the level sets approximate minimal surfaces, and the squared term involving $H$ and $\Delta u$ vanishes. The bound $A^{\prime}(t)\geq\int_{\Sigma_{t}}R_{\tilde{g}}/|\nabla u|\,d\sigma\geq 0$ remains valid.
@@
Quantitative preservation of sub-extremality factor. For strictly sub-extremal initial data with $A(0)>8\pi|J|$, define the sub-extremality factor:
$\mathcal{S}(t):=1-\frac{64\pi^{2}J^{2}}{A(t)^{2}}.$
Since $A(t)\geq A(0)$ and $J(t)=J$ is constant:
$\mathcal{S}(t)=1-\frac{64\pi^{2}J^{2}}{A(t)^{2}}\geq 1-\frac{64\pi^{2}J^{2}}{A(0)^{2}}=\mathcal{S}(0)>0.$
The sub-extremality factor is non-decreasing along the flow and remains strictly positive if it starts strictly positive. This ensures the monotonicity formula (Theorem 6.30) has a non-negative integrand throughout the flow.
@@
    Dain–Reiris inequality (Theorem 7.2): The proof of $A\geq 8\pi|J|$ in [22] relies on the stability condition through a variational argument. Unstable MOTS can violate this bound.
@@
Remark 7.7 (Bootstrap Structure of the Sub-Extremality Argument).
A careful reader may wonder about a potential circularity: the Dain–Reiris inequality is proven for MOTS, but the level sets $\Sigma_{t}$ for $t>0$ are not MOTS. How can we claim $A(t)\geq 8\pi|J|$ for all $t$?
Resolution: The argument is not a re-application of Dain–Reiris at each $t$. Instead:
  1. (1)
    Initial bound (Dain–Reiris): At $t=0$, the outermost MOTS $\Sigma_{0}=\Sigma$ satisfies $A(0)\geq 8\pi|J|$ by the Dain–Reiris theorem [22]. This is a one-time application at the MOTS only.
  2. (2)
    Angular momentum conservation: By Theorem 6.14, $J(t)=J$ is constant for all $t\in[0,1]$. This uses Stokes’ theorem and the vacuum condition in the exterior.
  3. (3)
    Area monotonicity (proven independently): By the AMO theory [1], the area $A(t)$ of level sets is non-decreasing: $A(t)\geq A(0)$. This follows from the $p$-harmonic structure and $R_{\tilde{g}}\geq 0$, and does not require $\Sigma_{t}$ to be a MOTS.
  4. (4)
    Conclusion (algebraic): Combining (1), (2), (3):
    $A(t)\geq A(0)\geq 8\pi|J|=8\pi|J(t)|.$
    No circularity exists because the Dain–Reiris inequality is used only at $t=0$, and the preservation for $t>0$ follows from the independent monotonicity of area.
Key point: The Dain–Reiris inequality and the AMO area monotonicity are logically independent theorems. Dain–Reiris applies to MOTS and uses MOTS-specific variational arguments. AMO area monotonicity applies to level sets of $p$-harmonic functions and uses the Bochner technique with $R_{\tilde{g}}\geq 0$. The sub-extremality preservation is the combination of these two independent results.
@@ 8 Synthesis: Complete Proof
Definition 8.1 (Cylindrical End Boundary Conditions—Formal Specification).
Let $(\bar{M},\bar{g})$ be the Jang manifold with cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$ at the MOTS. The cylindrical end boundary conditions for the AM-Lichnerowicz equation and AMO flow are:
AM-Lichnerowicz equation ($8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=\Lambda_{J}\phi^{-7}$):
  1. (BC-L1)
    Asymptotic Dirichlet at MOTS: $\phi(t,y)\to 1$ as $t\to\infty$, uniformly in $y\in\Sigma$.
  2. (BC-L2)
    Exponential decay: $|\phi(t,y)-1|\leq Ce^{-\kappa t}$ for some $\kappa=\kappa(\lambda_{1}(L_{\Sigma}))>0$.
  3. (BC-L3)
    Derived Neumann: $\partial_{t}\phi(t,y)\to 0$ as $t\to\infty$ (consequence of (BC-L2)).
  4. (BC-L4)
    Asymptotic Dirichlet at infinity: $\phi(x)\to 1$ as $r\to\infty$ in the asymptotically flat end.
AMO $p$-harmonic potential ($\Delta_{p}u=0$):
  1. (BC-A1)
    Dirichlet at MOTS: $u(t,y)\to 0$ as $t\to\infty$ along the cylindrical end.
  2. (BC-A2)
    Dirichlet at infinity: $u(x)\to 1$ as $r\to\infty$ in the asymptotically flat end.
  3. (BC-A3)
    Monotonicity: $u$ is strictly monotone from the MOTS (value 0) to infinity (value 1).
Well-posedness statements:
  • Conditions (BC-L1)–(BC-L4) determine a unique positive solution $\phi>0$ (Theorem 5.8).
  • Conditions (BC-A1)–(BC-A3) determine a unique solution $u$ for each $p\in(1,2]$.
  • The boundary conditions are asymptotic, not imposed on a compact boundary.
Remark 8.2 (Hypothesis Usage Summary).
The four hypotheses enter the proof as follows:
  • (H1)
    DEC: Ensures $R_{\tilde{g}}\geq 0$ after conformal transformation (Stage 2), which drives AMO monotonicity (Stage 6).
  • (H2)
    Axisymmetry: Defines angular momentum $J$, guarantees axisymmetric solutions at every stage, and enables $J$-conservation (Stage 4).
  • (H3)
    Exterior vacuum: Ensures Komar and ADM angular momenta coincide; enables clean asymptotics for boundary evaluation (Stage 7).
  • (H4)
    Strictly stable MOTS: Guarantees $|\theta^{-}|>0$, ensuring proper cylindrical blow-up and correct boundary values at $t=0$ (Stage 7).
Remark 8.3 (Strictly Stable vs. Marginally Stable MOTS).
A MOTS $\Sigma$ is characterized by its stability operator $L_{\Sigma}$, defined as the linearization of the null expansion $\theta^{+}$ under normal deformations:
$L_{\Sigma}\psi=-\Delta_{\Sigma}\psi+2\omega\cdot\nabla_{\Sigma}\psi+\left(\tfrac{1}{2}\mathrm{Sc}_{\Sigma}-|\chi^{+}|^{2}-|\omega|^{2}+\mathrm{div}_{\Sigma}\omega\right)\psi,$
where $\omega$ is the connection 1-form on the normal bundle and $\chi^{+}$ is the shear of the outgoing null normal.
Definition.
  • Strictly stable MOTS: The principal eigenvalue satisfies $\lambda_{1}(L_{\Sigma})>0$. This implies $|\theta^{-}|>0$ (the ingoing expansion is strictly negative).
  • Marginally stable MOTS: $\lambda_{1}(L_{\Sigma})=0$, with a non-trivial kernel. This corresponds to the “critical” case $\theta^{-}=0$ at isolated points or the entire surface.
  • Unstable MOTS: $\lambda_{1}(L_{\Sigma})<0$.
Role in the Proof. The strictly stable assumption $\lambda_{1}(L_{\Sigma})>0$ is used in the following essential ways:
  1. (1)
    Jang equation blow-up rate (Stage 1): The Jang solution $f$ satisfies $f\sim C_{0}\ln(1/s)$ near the MOTS, where $C_{0}=|\theta^{-}|/2$. For strictly stable MOTS, $|\theta^{-}|>0$ gives $C_{0}>0$, ensuring the Jang metric becomes a proper cylinder with metric $dt^{2}+g_{\Sigma}$. For marginally stable MOTS with $\theta^{-}=0$, the blow-up rate degenerates ($C_{0}=0$), and the Jang solution may have a different asymptotic structure (e.g., conical rather than cylindrical).
  2. (2)
    Spectral gap for elliptic regularity (Stages 2–3): The exponential decay $|\phi-1|=O(e^{-\kappa t})$ on the cylinder relies on the spectral gap $\kappa=\kappa(\lambda_{1})>0$. When $\lambda_{1}=0$, the decay becomes polynomial rather than exponential, requiring different function space techniques (cf. Mazzeo–Pacard [96]).
  3. (3)
    Boundary value extraction (Stage 7): The MOTS boundary value $m_{H,J}(0)$ is computed using $H_{\tilde{g}}(\Sigma)=0$ (the conformal metric has minimal boundary). This follows from $H_{\bar{g}}(\Sigma)=0$ (Lemma 8.4) and the conformal scaling $H_{\tilde{g}}=\phi^{-2}(H_{\bar{g}}+4\bar{\nu}(\ln\phi))$, where $\bar{\nu}(\ln\phi)\to 0$ by the exponential decay of $\phi-1$.
Marginally Stable Case: Status. The marginally stable case $\lambda_{1}(L_{\Sigma})=0$ (which includes degenerate horizons such as extremal Kerr cross-sections) requires substantial modifications:
  • The Jang manifold structure near the MOTS may differ from a cylinder.
  • Polynomial decay spaces (Melrose–Mendoza b-calculus) replace exponential decay spaces.
  • The boundary value analysis requires careful treatment of logarithmic terms.
This paper addresses only the strictly stable case. Extension to marginally stable MOTS is an important direction for future work (see Section 10).
@@
Let $(M,g,K)$ be asymptotically flat, axisymmetric data satisfying DEC with outermost stable MOTS $\Sigma$.
Stage 1: By Theorem 4.12, solve the axisymmetric Jang equation to obtain $(\bar{M},\bar{g})$ with cylindrical ends at $\Sigma$.
Stage 2: By Theorem 5.8, solve the AM-Lichnerowicz equation to obtain $\tilde{g}=\phi^{4}\bar{g}$ with $R_{\tilde{g}}\geq 0$.
@@
Stage 4: By Theorem 6.14, $J(t)=J$ for all $t\in[0,1]$.
Stage 5: By Theorem 7.2, $A(t)\geq 8\pi|J|$ for all $t$.
Stage 6: By Theorem 6.30, $m_{H,J}(t)$ is monotone increasing.
Stage 7: Boundary values as $p\to 1^{+}$.
@@
Step 1: Geometric setup on the Jang manifold. On the Jang manifold $(\bar{M},\bar{g})$, the MOTS $\Sigma$ becomes the boundary of the cylindrical end. The key property is that the mean curvature $H_{\bar{g}}$ of $\Sigma$ in $(\bar{M},\bar{g})$ vanishes, i.e., $\Sigma$ is a minimal surface in the Jang metric. We now prove this fact.
Important clarification: The physical MOTS condition is $\theta^{+}=H_{g}+\mathrm{tr}_{\Sigma}K=0$, where $H_{g}$ is the mean curvature in the physical metric $g$. This does not imply $H_{g}=0$; rather, for non-time-symmetric data with $K\neq 0$, we have $H_{g}=-\mathrm{tr}_{\Sigma}K\neq 0$ generically. The property $H_{\bar{g}}=0$ (mean curvature in the Jang metric) is a separate geometric fact that follows from the cylindrical end structure of the Jang solution, as we now derive.
Detailed derivation of $H_{\bar{g}}|_{\Sigma}=0$: The Jang surface $\Gamma_{f}=\{(x,f(x)):x\in M\}$ is embedded in $(M\times\mathbb{R},g+dt^{2})$. Near the MOTS $\Sigma$, the Jang solution $f$ blows up as (see Theorem 4.12(iii)):
@@
$\bar{g}=g+df\otimes df=g+\frac{C_{0}^{2}ds\otimes ds}{s^{2}}+O(s^{-1}).$
In the cylindrical coordinate $t=C_{0}\ln(1/s)$ (so $s=e^{-t/C_{0}}$), this becomes:
@@
Rigorous proof of $H_{\bar{g}}|_{\Sigma}=0$: Consider the slice $\Sigma_{t}:=\{t\}\times\Sigma$ in the cylindrical end. The second fundamental form of $\Sigma_{t}$ in $(\bar{M},\bar{g})$ is computed from the Lie derivative of the metric:
$h_{ij}(t)=\frac{1}{2}(\mathcal{L}_{\partial_{t}}\bar{g})_{ij}|_{\Sigma_{t}}=\frac{1}{2}\partial_{t}(g_{\Sigma_{t}})_{ij}.$
By Theorem 4.12(iii), the metric on $\Sigma_{t}$ converges exponentially to the MOTS metric:
$g_{\Sigma_{t}}=g_{\Sigma}+O(e^{-\beta_{0}t}),\quad\partial_{t}g_{\Sigma_{t}}=O(e^{-\beta_{0}t}).$
Therefore the second fundamental form satisfies $h_{ij}(t)=O(e^{-\beta_{0}t})$, and the mean curvature:
$H_{\bar{g}}(\Sigma_{t})=\mathrm{tr}_{g_{\Sigma_{t}}}h(t)=O(e^{-\beta_{0}t})\to 0\quad\text{as }t\to\infty.$
@@
$H_{\bar{g}}|_{\Sigma}:=\lim_{t\to\infty}H_{\bar{g}}(\Sigma_{t})=0.$ (93)
This is the key geometric fact: the MOTS $\Sigma$ is a minimal surface in the Jang metric $\bar{g}$.
Physical interpretation: The vanishing $H_{\bar{g}}|_{\Sigma}=0$ does not follow directly from the MOTS condition $\theta^{+}=0$. Rather, it follows from the cylindrical structure of the Jang blow-up: near the MOTS, the Jang manifold asymptotes to an infinite cylinder, and cross-sections of product cylinders are totally geodesic (hence have zero mean curvature).
@@
Remark 8.5 (Boundary Conditions on the MOTS).
The Dirichlet condition $\phi|_{\Sigma}=1$ and the Neumann condition $\partial_{\nu}\phi|_{\Sigma}=0$ require careful justification, as the MOTS $\Sigma$ is not a boundary in the classical PDE sense but rather the “end” of a cylindrical neck.
Why $\phi|_{\Sigma}=1$ (Dirichlet):
  • The MOTS $\Sigma$ becomes the boundary of the cylindrical end in the Jang manifold. In cylindrical coordinates $(t,y)\in[0,\infty)\times\Sigma$, $\Sigma$” corresponds to $t\to\infty$.
  • The AM-Lichnerowicz equation is solved on the compact region $\bar{M}_{T}=\bar{M}\setminus\{t>T\}$ with Dirichlet data $\phi=1$ on the “cap” $\{t=T\}\times\Sigma$.
  • As $T\to\infty$, the solutions converge to the unique solution on the full $\bar{M}$ with asymptotic value $\phi\to 1$ along the cylindrical end.
  • This is not an arbitrary choice: the normalization $\phi=1$ at the MOTS ensures $A_{\tilde{g}}(\Sigma)=A_{g}(\Sigma)$, which is essential for the correct boundary value of $m_{H,J}(0)$.
Why $\partial_{\nu}\phi|_{\Sigma}=0$ (Neumann):
  • The exponential decay of $\phi-1$ on the cylinder (from the spectral gap of $\Delta_{\Sigma}$) implies all derivatives also decay: $|\partial_{t}^{k}(\phi-1)|=O(e^{-\kappa t})$.
  • The limiting value $\partial_{\nu}\phi|_{\Sigma}=\lim_{t\to\infty}\partial_{t}\phi=0$ is a consequence of the equation, not an imposed condition.
  • Physically, this reflects the fact that $\Sigma$ is a “barrier” for the conformal factor: the geometry cannot “leak” mass through the minimal surface.
Summary of well-posedness: The AM-Lichnerowicz equation on $\bar{M}$ with:
  1. (i)
    Asymptotic condition $\phi\to 1$ at spatial infinity, and
  2. (ii)
    Asymptotic condition $\phi\to 1$ along the cylindrical end at $\Sigma$
admits a unique positive solution $\phi>0$. Both conditions are asymptotic behavior constraints, not boundary data in the classical sense. The Neumann condition $\partial_{\nu}\phi|_{\Sigma}=0$ is derived, not imposed.
Rigorous justification of $\partial_{\nu}\phi|_{\Sigma}=0$ via elliptic estimates: We provide three independent arguments for completeness:
  1. (i)
    Variational argument: The AM-Lichnerowicz equation is the Euler–Lagrange equation for the functional $\mathcal{E}[\phi]=\int 8|\nabla\phi|^{2}+R_{\bar{g}}\phi^{2}+\frac{\Lambda_{J}}{6}\phi^{-6}$. On a manifold with minimal boundary (which $\Sigma$ is, by $H_{\bar{g}}|_{\Sigma}=0$), the natural boundary condition for critical points is Neumann: $\partial_{\nu}\phi=0$ (see [42, Proposition 3.2]).
  2. (ii)
    Exponential decay argument: On the cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$, write $\phi=1+\psi$ where $\psi$ solves a linear elliptic equation with source terms decaying as $O(e^{-\beta_{0}t})$. By standard elliptic estimates on cylinders (Lockhart–McOwen theory), $\psi$ and all its derivatives decay exponentially: $|\partial_{t}^{k}\partial_{\Sigma}^{\ell}\psi|=O(e^{-\kappa t})$ for some $\kappa>0$ determined by the spectral gap. In particular, $\partial_{t}\phi=\partial_{t}\psi=O(e^{-\kappa t})\to 0$ as $t\to\infty$.
  3. (iii)
    Uniqueness argument: The AM-Lichnerowicz equation on $\bar{M}$ with boundary $\partial\bar{M}=\Sigma$ (at infinity along the cylinder) and asymptotic condition $\phi\to 1$ at spatial infinity admits a unique solution. This solution is obtained as the limit of Dirichlet problems on $\bar{M}_{T}=\bar{M}\setminus(\{t>T\}\times\Sigma)$ with $\phi|_{\{t=T\}\times\Sigma}=1$. By the maximum principle, $\phi\leq 1$ throughout (since $\Lambda_{J}\geq 0$ makes 1 a supersolution). The limit $T\to\infty$ converges to the unique solution with $\phi|_{\Sigma}=1$ and (by the exponential decay of gradients) $\partial_{\nu}\phi|_{\Sigma}=0$.
@@
$f(s,y)=C_{0}\ln(1/s)+\mathcal{A}(y)+O(s^{\alpha})=C_{0}t+\mathcal{A}(y)+O(e^{-\alpha t}).$
@@
Rigorous justification via induced metric formula. Let $\{e_{1},e_{2}\}$ be an orthonormal frame for $T\Sigma$ in the metric $g$. The induced metric components on $\Sigma$ are:
$(\bar{g}|_{\Sigma})_{ab}=\bar{g}(e_{a},e_{b})=g(e_{a},e_{b})+df(e_{a})\cdot df(e_{b}).$
Since $f$ blows up in the normal direction with $\nabla_{g}f=C_{0}\nu/s+O(1)$ (where $\nu\perp T\Sigma$), we have:
$df(e_{a})=g(\nabla f,e_{a})=C_{0}s^{-1}g(\nu,e_{a})+O(1)=0+O(1)$
because $\nu\perp e_{a}$. Thus $df(e_{a})=O(1)$ remains bounded, and in the limit $s\to 0$:
$\lim_{s\to 0}(\bar{g}|_{\Sigma})_{ab}=g(e_{a},e_{b})+\lim_{s\to 0}O(1)\cdot O(1)=(g|_{\Sigma})_{ab}.$
More precisely, on the slices $\Sigma_{T}$ at cylindrical height $T$, the tangential gradient $|df^{\text{tan}}|$ decays as $O(e^{-\beta_{0}T})$, so $|(\bar{g}-g)|_{\Sigma_{T}}|=O(e^{-2\beta_{0}T})\to 0$.
@@
Remark 8.7 (Summary of MOTS Boundary Value Proof).
The key equation $m_{H,J}(0)=\sqrt{A/(16\pi)+4\pi J^{2}/A}$ follows from:
  1. (1)
    Minimality in $\bar{g}$: The MOTS $\Sigma$ is the boundary of the cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$ in the Jang manifold. Cylindrical slices are asymptotically totally geodesic, so $H_{\bar{g}}|_{\Sigma}=0$.
  2. (2)
    Neumann boundary for $\phi$: On the cylinder, exponential decay $\phi=1+O(e^{-\kappa t})$ implies $\partial_{\nu}\phi|_{\Sigma}=0$. This plus $\phi|_{\Sigma}=1$ yields $H_{\tilde{g}}|_{\Sigma}=\phi^{-2}(H_{\bar{g}}+4\phi^{-1}\partial_{\nu}\phi)|_{\Sigma}=0$.
  3. (3)
    Hawking mass of minimal surface: For $H_{\tilde{g}}|_{\Sigma}=0$: $m_{H}(\Sigma)=\sqrt{A_{\tilde{g}}(\Sigma)/(16\pi)}$.
  4. (4)
    Area preservation: Since $df|_{\Sigma}$ is purely normal and $\phi|_{\Sigma}=1$: $A_{\tilde{g}}(\Sigma)=A_{\bar{g}}(\Sigma)=A_{g}(\Sigma)=A$.
  5. (5)
    Conclusion: $m_{H,J}(0)=\sqrt{m_{H}^{2}+\frac{4\pi J^{2}}{A}}=\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}}$.
This is an equality, not merely a lower bound.
@@
    Limit of $m_{H,J}(t)$: The value $m_{H,J}(0)$ is defined as $\lim_{t\to 0^{+}}m_{H,J}(t)$. By the continuity of area and the fact that $\Sigma_{t}\to\Sigma$ in the Hausdorff topology (with controlled curvature from the $p$-harmonic structure), this limit equals the AM-Hawking mass computed directly on $\Sigma$ via Lemmas 8.4 and 8.6.
@@
    Mean curvature regularity: Since the level sets $\Sigma_{t}$ are $C^{1,\beta}$ regular for $p\in(1,2]$ [1], the mean curvature $H$ and second fundamental form $h$ are well-defined almost everywhere. The Hawking mass integral $\int_{\Sigma_{t}}H^{2}\,dA$ is finite for regular level sets.
@@
    $m_{H}(t)$ $=\sqrt{\frac{A(t)}{16\pi}}\left(1-W(t)\right)$
    $\approx\frac{r(t)}{2}\left(\frac{2M_{\mathrm{ADM}}}{r(t)}+O(r(t)^{-1-\tau})\right)$
@@
    The expansion uses the standard ADM mass formula: for coordinate spheres $S_{r}$, $\int_{S_{r}}H^{2}dA=16\pi-32\pi M_{\mathrm{ADM}}/r+O(r^{-1-\tau})$, giving $1-W(t)=2M_{\mathrm{ADM}}/r(t)+O(r^{-1-\tau})$. ∎
@@
Conclusion: By the monotonicity from Stage 6 and the mass chain from Lemma 5.9:
@@
The last inequality uses the lower bound analysis from Stage 7 at the MOTS, which becomes an equality for Kerr initial data. ∎